### v1

# Whose Boat Does it Float? Improving Personalization in Preference Tuning via Inferred User Personas

Abstract
LLMs are tuned to follow instructions (aligned) by learning which of two outputs users prefer for a prompt. However, this preference data format does not convey why users prefer responses that are chosen or rejected, so LLMs trained on these datasets cannot tailor responses to varied user needs. To surface these parameters of personalization, we apply abductive reasoning to preference data, inferring needs and interests of users, i.e. personas, that may prefer each output. We test this idea in two steps: Persona Inference (PI)—abductively inferring personas of users who prefer chosen or rejected outputs—and Persona Tailoring (PT)—training models to tailor responses to personas from PI. We find: 1) LLMs infer personas accurately explaining why different users may prefer both chosen or rejected outputs; 2) Training on preference data augmented with PI personas via PT boosts personalization, enabling models to support user-written personas; and 3) Rejected response personas form harder personalization evaluations, showing PT better aids users with uncommon preferences versus typical alignment methods. We argue for an abductive view of preferences for personalization, asking not only which response is better but when, why, and for whom. [^1] Code/data are publicly available at https://github.com/Pinafore/alignment-personalization
1 Every Preference Happens for a Reason
Current methods for aligning large language models (LLMs) predominantly rely on preference data Ji et al. (2023), which is created by finding: for an input prompt, which of two outputs do users prefer? LLMs are then trained to give responses like the majority chosen response and unlike the rejected one, improving LLM output quality in dialogue Köpf et al. (2024), question answering (Fan et al., 2019, QA), and summarization Völske et al. (2017).
While useful, preference datasets assume chosen responses are universally better, failing to consider why users prefer responses. In reality, some users may prefer rejected outputs, even if their reasons are less common. In Figure 2 (left), most users prefer the chosen output for taking brownies to a cake sale. While users valuing simplicity may prefer this direct response, practical users may prefer the rejected one which also has packaging logistics (Figure 2, middle). As users prefer responses for varied reasons Kirk et al. (2024b), models tuned on preference data should personalize outputs to meet these specific, individual needs Salemi et al. (2024).
A common, interpretable method for personalization involves users verbalizing their needs and interests as an extra input (e.g. “I like short answers”)—forming a system prompt persona to tailor model responses Zhang et al. (2024c). However, current preference data formats do not have personas conveying why users prefer responses, omitting signals for training personalized models Lee et al. (2024).
Figure 1: Alignment methods with typical preference datasets like DPO cannot fully cater to a user’s specified personas. To overcome this, we train models on preference data augmented with LLM-inferred personas, which we call persona tailoring.
Figure 2: Overview of this paper. Preference data has a prompt, chosen response, and rejected response (left). Most users prefer the chosen response, but there are valid reasons and personas of users that may prefer either response, which we uncover via abductive reasoning in Persona Inference (PI, middle). We then study Persona Tailoring (PT) to use the personas for data augmentation in few-shot prompts, fine-tuning, and direct preference optimization—enhancing personalization (right).
The only way for models tuned on existing preference data formats to use a persona is by adding it to the prompt in inference, hoping the model treats it as an extra instruction to follow Deb et al. (2022). However, typical preference tuning methods like direct preference optimization (Rafailov et al., 2024, DPO) lack this personalization. In Figure 1, DPO tells a sober user to hire a bartender (left) for their party, while a user desiring a small list gets 10 ideas (right); such responses are unhelpful, inappropriate, and fail to support the users’ specified personas.
As models’ poor personalization stems from a lack of training on personas, we propose abductive reasoning Peirce (1974) to augment preference training data with LLM-inferred personas. Abduction infers hidden contexts that explain a given outcome Zhao et al. (2024a). Similarly, we adopt this reasoning to infer (i.e. adduce) hidden needs, interests, and traits of users, i.e., personas Tseng et al. (2024), that explain why chosen and rejected outputs may be preferred. If LLMs surface valid personas of users who may prefer chosen or rejected responses, we can attach them to preference datasets and train models that tailor to user-specified needs, enhancing model personalization (Figure 1, blue).
We segment this idea into a two-task sequence: 1) Persona Inference (PI, section 2): adducing personas for preference data responses (Figure 2, middle); and 2) Persona Tailoring (PT, section 4): producing outputs tailored to personas from PI (Figure 2, right).
We first thoroughly test LLM abilities in PI on dialogue Bai et al. (2022a), QA Ji et al. (2024), and education Balepur et al. (2024b) preference data. LLM personas accurately explain why users may prefer chosen and rejected outputs; LLaMA-405B has 91% accuracy, judged by GPT-4o (section 3.1). Also, GPT-4o often judges chosen and rejected response personas as tied in quality (section 3.2), and qualitatively, rejected ones are rated as highly-plausible but applying to less users (section 3.3). Thus, users may prefer rejected outputs for uncommon but valid reasons. Personas also describe implicit differences Hoyle et al. (2023) in chosen and rejected responses; we find in BeaverTails Ji et al. (2024), users in chosen response personas are “meticulous”, while rejected ones are “direct”, showing the data annotators tend to prefer verbose responses Zheng et al. (2024).
As PI yields accurate personas, we augment preference data with PI personas using LLaMA-405B. We then train LLaMA-8B on this data for the reverse task of PT: using prompts and inferred personas as inputs to give tailored responses (Figure 2, right). We test three strategies: prompting (Brown et al., 2020, $\textsc{PT}_{\textsc{fs}}$), fine-tuning (Chung et al., 2024, $\textsc{PT}_{\textsc{sft}}$), and DPO (Rafailov et al., 2024, $\textsc{PT}_{\textsc{dpo}}$).
Training on LLM personas via PT largely aids personalization (section 5.1), with $\textsc{PT}_{\textsc{dpo}}$ as the strongest method (section 5.2). Further, DPO tuned on preference datasets without personas cannot always tailor to personas in inference; notably, $\textsc{PT}_{\textsc{dpo}}$ is judged as much stronger than DPO on uncommon but valid needs in the personas from rejected responses (58% average improvement in personalization), showing rejected responses form valuable, harder evaluations for personalization (section 5.3). Lastly, we have ten users write diverse personas and qualitatively assess $\textsc{PT}_{\textsc{dpo}}$ and DPO outputs (section 5.4). The same users find $\textsc{PT}_{\textsc{dpo}}$ better personalizes to their written needs, showing its utility to aid users downstream.
We argue for an abductive view of preferences, capturing not only which outputs users prefer but which users and for what reasons. In doing so, we can find valid user needs that may be overlooked in rejected responses (section 3) and ensure LLMs support them (section 5.3). It also improves personalization (section 5), such as augmentation via PI. Our contributions are:
1) We study abductive reasoning on preference data via persona inference (PI) to show LLMs can infer why users may prefer chosen and rejected outputs.
2) We employ PI for data augmentation and release persona-augmented versions of question answering, dialogue, and education preference datasets.
3) We show persona tailoring—training on persona-augmented preference data—enhances personalization for LLM-inferred and user-specified needs.
2 Persona Inference Setup
The first step in our personalization method is inferring why users may prefer each output in standard preference data (Figure 2, middle). We adopt abduction Peirce (1974)—which explains given outcomes—to infer why preference data responses may be preferred via Persona Inference (PI):
$\bullet\;\textsc{PI}(p,r_{1},r_{2})\rightarrow\mathcal{P}_{1}$: For prompt $p$ and responses $r_{1}$ and $r_{2}$, the LLM gives a persona $\mathcal{P}_{1}$ such that a user described by $\mathcal{P}_{1}$ would prefer $r_{1}$ over $r_{2}$. If $r_{1}$ is the chosen response and $r_{2}$ is the rejected one, $\mathcal{P}_{1}$ will describe a user preferring the chosen response, and vice versa if $r_{1}$ is the rejected response.
Following the taxonomy of Chen et al. (2024a), our personas are closest to demographic personas. However, we infer high-level traits (Figure 2, middle), like information needs, interests, and personality types, that alter how users want outputs shown, rather than protected attributes (e.g. race) to avoid stereotyping Kantharuban et al. (2024). This section gives our PI models (section 2.1) and datasets (section 2.2).
2.1 Models
We test nine LLMs in PI: Claude (Anthropic, 2023, Sonnet, Haiku, Opus), GPT (Achiam et al., 2023, 3.5, 4, 4o), and LLaMA-3.1 Instruct (Dubey et al., 2024, 8B, 70B, 405B). We use 0 temperature and five-shot prompts written by one author, using rules in Schulhoff et al. (2024). Exemplars use the format below, where highlights show generated text:
Prompt 2.1: Few-Shot Persona Inference Prompt Prompt: $p$ Chosen Response: $r_{1}$ Rejected Response: $r_{2}$ Persona: $\mathcal{P}_{1}$
We append text to Prompt 2.1 asking for a “short, one-sentence description of the user’s preference”. We also specify the persona must have “high-level characteristics” to avoid stereotypes and not have exact phrases in the prompts or responses to avoid the trivial solution of repeating $r_{1}$ for $\mathcal{P}_{1}$ (section 3.3).
2.2 Datasets
PI requires preference data with an input prompt $p$ and responses $\mathcal{R}=[r_{1},r_{2}]$. For a thorough evaluation, we use four datasets in question answering (QA), dialogue, and education—three domains that benefit from personalization Zhang et al. (2024c):
1) BeaverTails Ji et al. (2024) has advice questions $p$ and candidate answers $\mathcal{R}$ on 14 harm categories including politics, privacy, and unethical behavior.
2) Stanford Human Preferences (Ethayarajh et al., 2022, SHP) has Reddit post questions $p$ on r/ask and advice forums with user-written answers $\mathcal{R}$.
3) Anthropic HHH Bai et al. (2022a) has human inputs $p$ and assistant responses $\mathcal{R}$ from real dialogues. We use single-turn dialogues for simplicity.
4) Mnemonic Balepur et al. (2024b) has vocabulary terms $p$ and keyword mnemonics $\mathcal{R}$. Mnemonics are study aids that help users learn $p$’s meaning.
The datasets have users rate the “better” response $r\in\mathcal{R}$, where “better” means more helpful/harmless for (1) and (3), gets more Reddit upvotes for (2), and aids learning in (4). The overall better $r$ is chosen ($r_{\texttt{C}}$) and the other is rejected ($r_{\texttt{R}}$). For each entry, we alter inputs $r_{1}$ and $r_{2}$ in PI to get chosen personas $\mathcal{P}_{\texttt{C}}=\textsc{PI}(p,r_{1}=r_{\texttt{C}},r_{2}=r_{\texttt{R}})$ and rejected personas [^2] We use “rejected” for brevity, not to imply they are worse. $\mathcal{P}_{\texttt{R}}=\textsc{PI}(p,r_{1}=r_{\texttt{R}},r_{2}=r_{\texttt{C}})$ for $r_{\texttt{C}}$ and $r_{\texttt{R}}$. As $r_{\texttt{R}}$ is preferred less often, we assume $\mathcal{P}_{\texttt{R}}$ has less common/popular needs (tested in section 3.3).
In BeaverTails and SHP, some responses $r_{\texttt{R}}$ are deliberately low-quality, with harmful or inaccurate content Liu et al. (2024b). These are out-of-scope, as all $\mathcal{P}_{\texttt{R}}$ will be undesirable [^3] Such as “the user is unethical” or “likes misleading users”. and we do not want to have models tailor to them (section 9). Thus, we use a split of outputs labeled “safe” in BeaverTails and outputs with 10+ upvotes in SHP. Post-filtering, we randomly sample 300 entries in each dataset to form 600 PI inputs (data details in Appendix A.1).
3 Understanding LLM-Inferred Personas
Figure 3: GPT-4o judgments on if LLM personas accurately infer users who prefer chosen/rejected responses. Personas are highly accurate and chosen/rejected persona accuracy gaps are small, so users may prefer rejected outputs for valid reasons.
Figure 4: Persona quality comparison. Excluding BeaverTails, GPT-4o rates chosen and rejected personas as similar in quality.
Before directly using LLM personas to train personalized models (section 4), we first thoroughly assess their quality and differences in chosen and rejected personas, providing insights into the types of personas LLMs infer. This section finds LLMs’ chosen and rejected personas: 1) accurately explain why users may prefer each preference dataset response (section 3.1); 2) convey similarly-valid needs (section 3.2, section 3.3); and 3) reveal preference data trends (section 3.4), showing they are reliable signals for augmenting preference data.
3.1 LLMs Accurately Infer Personas
We first ensure LLM personas satisfy abduction’s goal: justifying why its input response is better (section 2). To test this, we have GPT-4o, a strong judge Kim et al. (2024), pick the best response out of $r_{\texttt{C}}$ and $r_{\texttt{R}}$ for a prompt and persona. A chosen persona $\mathcal{P_{\texttt{C}}}$ is accurate if GPT-4o picks $r_{\texttt{C}}$ over $r_{\texttt{R}}$ (same for $\mathcal{P_{\texttt{R}}}$). GPT-4o shows 86% agreement with one Ph.D. student (Appendix A.4), so the metric is reliable.
LLMs infer accurate $\mathcal{P_{\texttt{C}}}$ and $\mathcal{P_{\texttt{R}}}$; the judge picks the intended response in most cases (Figure 3). $\mathcal{P_{\texttt{R}}}$ is always less accurate, consistent with work that shows LLMs struggle to justify incorrect answers Michael et al. (2023); Balepur et al. (2024a). However, some models have small accuracy gaps (0.06 for L-405B), so while $\mathcal{P_{\texttt{R}}}$ is harder to infer, it can still identify plausible needs of users. Finally, Mnemonic accuracy is lowest, as LLMs must infer why outputs aid learning, which is likely harder than why they are helpful or harmless. Thus, in specific, subjective domains Padmakumar et al. (2024), it may be useful to directly elicit why annotators prefer responses in preference data collection, rather than relying on LLMs to infer them.
3.2 LLMs Judge Personas as Similar Quality
As LLMs infer accurate chosen/rejected personas $\mathcal{P}_{\texttt{C}}$/$\mathcal{P}_{\texttt{R}}$, we now compare $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$’s quality. If they are judged as similar-quality, we can be more confident that $\mathcal{P}_{\texttt{R}}$ has needs as valid as $\mathcal{P}_{\texttt{C}}$. To do this, we 0-shot prompt GPT-4o to judge if $\mathcal{P}_{\texttt{C}}$ or $\mathcal{P}_{\texttt{R}}$ is higher-quality, yielding a persona preference $y$. We shuffle the personas and set $y=\texttt{C}\ \textrm{(}\text{or }\texttt{R}$) if $\mathcal{P}_{\texttt{C}}$ (or $\mathcal{P}_{\texttt{R}}$) win in both orders; otherwise, $y=\texttt{Tie}$.
On all datasets except BeaverTails, GPT-4o rates $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ as similar in quality (Figure 4); the two persona types have very similar win rates (mean difference of 0.1), further suggesting users can prefer rejected outputs for reasons as valid as chosen ones. The large win rate difference of 0.3 on BeaverTails implies $\mathcal{P}_{\texttt{C}}$ conveys user needs that GPT-4o tends to prefer over the needs in $\mathcal{P}_{\texttt{R}}$; we find these preferences may be linked to verbose responses in section 3.4.
3.3 Humans Rate Personas as High-Quality
To confirm LLM personas are high-quality beyond LLM judgments, one Ph.D. student rates 80 total chosen and rejected personas $\mathcal{P}$ from GPT-4o and L-405B on BeaverTails (randomly shuffled) across four axes. First, they rate Applicability from 1-5: how many users are expected to be represented by $\mathcal{P}$—a proxy for the popularity of reasons found in $\mathcal{P}$. Next, as a sanity check, the student gives binary labels for Plausibility: is the user in $\mathcal{P}$ likely to exist; Harmfulness: if $\mathcal{P}$ describes harmful or unethical traits we do not want to train models on; and Overfitting: if $\mathcal{P}$ directly repeats text in prompts or responses, rather than inferring high-level traits.
Figure 5: Qualitative sanity check of personas. Chosen and rejected personas are similarly plausible, harmless, and do not overfit, but rejected personas are considered less applicable.
Chosen and rejected personas are both plausible users, but rejected personas are less applicable (Figure 5), aligning with our intuition that rejected personas are less popular/common but nevertheless valid user needs (section 2.2). Chosen personas are also rarely harmful or overfit, indicating they can form high-quality personalization training signals (section 4).
Dataset Type Words (Saliency) Example Prompt Example Response Example Persona
BeaverTails Chosen multiple (0.94) What are the laws regarding animal cruelty in the United States? The United States has both state and federal laws that prohibit animal cruelty… The user prefers comprehensive, detailed information that covers the topic from multiple angles
meticulous (0.92)
diverse (0.92)
BeaverTails Rejected to-the-point (1.0) Where can I find legal advice regarding injury prevention? The CDC provides comprehensive legal advice and information on their website: cdc.gov/injury. The user is direct and to-the-point, preferring concise and specific information sources.
directness (0.90)
concise (0.90)
Mnemonic Chosen indirect (0.81) Ascertain Break it down to "assure + certain"; to ascertain is to make sure of something. The user is a logical thinker and prefers clear, step-by-step breakdowns to understand new concepts.
step-by-step (0.76)
essence (0.75)
Mnemonic Rejected strong (0.91) Zephyr Zephyr sounds like "sea-fear," like a gentle breeze that calms the fear of sailors asea. The user is a romantic thinker and prefers poetic, emotive associations.
story-like (0.90)
romantic (0.79)
Table 1: Persona word saliency. Running PI on preference data uncovers implicit differences in Chosen and Rejected responses.
3.4 Personas Help Describe Preference Data
We mainly use LLM personas for personalization (section 4), but they can also help describe trends in preference datasets. User preferences are implicit, but if LLMs articulate reasons behind these preferences, we can understand differences between chosen and rejected outputs Hoyle et al. (2023). To do this, we first form sets $\mathcal{S}_{\texttt{C}}$ and $\mathcal{S}_{\texttt{R}}$ with chosen and rejected personas from our three best LLMs: L-405B, Opus, and GPT-4o. We then find salient words in each set via $\mathcal{P}(w\,|\,\mathcal{S})$: the probability a word $w\in\mathcal{S}_{\texttt{C}}\cup\mathcal{S}_{\texttt{R}}$ appears given $w$ is in the chosen or rejected set $\mathcal{S}$.
Table 1 has words with the highest $\mathcal{P}(w\,|\,\mathcal{S})$ appearing 10+ times. [^4] Low-frequency words have inflated saliency scores (e.g. appearing only once in a rejected output yields a score of one). On BeaverTails, annotators choose detailed outputs (“multiple”, “meticulous”) and reject shorter ones (“to-the-point”, “concise”), explaining verbosity bias Zheng et al. (2024). On Mnemonic, learners mostly prefer mnemonics with breakdowns (“step-by-step”) that are not whimsical (“story-like”, “romantic”), helping educators write study aids that appeal to most learners. Thus, PI is also a useful content analysis tool for preference data, informing model designers and practitioners.
4 Persona Tailoring Setup
Personas from PI are accurate (section 3), so we use them to train more personalized models that use personas and prompts as inputs to produce tailored responses. We run PI with L-405B, our best open weight LLM, to add personas to preference data, and train L-8B on this new data for Persona Tailoring (PT):
$\bullet\;\textsc{PT}(p,\mathcal{P})\rightarrow r$: For prompt $p$ and persona $\mathcal{P}$, the LLM gives response $r$ for $p$ that is tailored to $\mathcal{P}$.
Our one-time augmentation strategy is similar to knowledge distillation Gou et al. (2021): a larger teacher LLM boosts personalization in a smaller student LLM, improving efficiency for long-term deployment. This section gives our datasets (section 4.1), strategies (section 4.2) and metrics (section 4.3) used for PT.
4.1 Datasets
PT needs persona-augmented preference datasets $\mathcal{D}_{\mathcal{P}}$ with a prompt $p$, responses [$r_{\texttt{C}}$, $r_{\texttt{R}}$], and personas [$\mathcal{P}_{\texttt{C}}$, $\mathcal{P}_{\texttt{R}}$]. We base $\mathcal{D}_{\mathcal{P}}$ on the BeaverTails, Anthropic HHH, and Mnemonic datasets $\mathcal{D}$ used in PI. [^5] We omit SHP since BeaverTails is also a QA dataset. We sample 2449, 1059, and 328 training entries from each $\mathcal{D}$ via section 2.2, with all splits in Appendix A.1. We sample 500 entries as BeaverTails and Anthropic test sets. Mnemonic is small, so we use a test set of 500 terms from its authors, which only has input prompts. We run PI with L-405B via section 2 on all splits except the Mnemonic test set to get personas $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ for each entry, yielding $\mathcal{D}_{\mathcal{P}}$.
BeaverTails Anthropic HHH Mnemonic
$\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
FS $\textsc{PT}_{\textsc{fs}}$+$\mathcal{P}_{\texttt{retr}}$ 62.5/17.2/20.2 60.7/14.2/25.1 +46.3 47.5/20.9/31.6 41.8/15.6/42.6 +9.6 44.3/28.5/27.2 46.4/20.5/33.1 +20.3
$\textsc{PT}_{\textsc{fs}}$+$\mathcal{P}_{\texttt{gold}}$ 68.7/14.5/16.9 62.9/15.9/21.3 +55.0 57.3/20.2/22.5 51.5/15.6/32.8 +32.9
SFT $\textsc{PT}_{\textsc{ft}}$+$\mathcal{P}_{\texttt{retr}}$ 44.6/31.7/23.7 33.5/28.6/37.8 +12.3 52.3/28.9/18.8 35.2/24.2/40.6 +20.0 40.8/38.3/20.9 35.2/35.2/29.5 +20.5
$\textsc{PT}_{\textsc{sft}}$+$\mathcal{P}_{\texttt{gold}}$ 46.7/32.0/21.2 38.2/29.6/32.2 +23.0 62.2/23.7/14.1 41.7/29.5/28.8 +40.6
DPO $\textsc{PT}_{\textsc{dpo}}$+$\mathcal{P}_{\texttt{retr}}$ 72.1/18.2/9.6 36.7/24.4/38.9 +36.8 54.1/26.1/19.8 21.0/23.7/55.3 +0.7 64.4/26.0/9.6 27.8/33.2/39.0 +28.6
$\textsc{PT}_{\textsc{dpo}}$+$\mathcal{P}_{\texttt{gold}}$ 66.3/21.4/12.2 40.9/28.5/30.7 +41.6 50.8/30.5/18.6 28.8/25.8/45.4 +12.0
Table 2: Win, tie, and loss rates of generation methods (FS, SFT, DPO) with and without personas $\mathcal{P}$ in pairwise comparisons from the Prometheus judge. Models that use personas often largely improve personalization without sacrificing response quality.
In test sets, using a persona from gold outputs $r_{\texttt{C}}$ or $r_{\texttt{R}}$ may be unrealistic, as it can leak signals in $r_{\texttt{C}}$ and $r_{\texttt{R}}$. Thus, for a test set prompt $p$, we retrieve a training example $\mathcal{E}$ with the most similar prompt Wang et al. (2022) via ColBERT Santhanam et al. (2022), and use personas from $\mathcal{E}$. We refer to personas from $\mathcal{E}$ as $\mathcal{P}_{\texttt{retr}}$, and from $r_{\texttt{C}}$/$r_{\texttt{R}}$ as $\mathcal{P}_{\texttt{gold}}$; we use both for thorough testing (examples in Appendix A.1). LLM personas are imperfect proxies for real needs, so users also write personas in section 5.4.
4.2 Personalization Techniques
We test three standard generation methods for PT:
1) Few-shot (Brown et al., 2020, FS) has 5 examples to produce $r_{\texttt{C}}$ from $p$ and $\mathcal{P}_{\texttt{C}}$ with the template:
Prompt 4.1: Few-Shot Persona Tailoring Prompt Prompt: $p$ Persona: $\mathcal{P}_{\texttt{C}}$ Response: $r_{\texttt{C}}$
2) Supervised fine-tuning (Chung et al., 2024, SFT) trains an LLM to generate $r_{\texttt{C}}$ from $p$ and $\mathcal{P}_{\texttt{C}}$ via the cross-entropy loss of next-token prediction:
${\mathcal{L}=\sum_{j=1}^{|r_{\texttt{C}}|}\log\textrm{P}(r_{j}\,|\,r_{1},...,r_{j-1},\langle p\cdot\mathcal{P}_{\texttt{C}}\rangle).}$
3) Direct preference optimization (Rafailov et al., 2024, DPO) tunes a model $\pi$ from the SFT model $\pi_{0}$, using $p$ and $\mathcal{P}_{\texttt{C}}$ as inputs $x$. DPO increases the probability gap for output sequences $r_{\texttt{C}}$ and $r_{\texttt{R}}$:
${\mathcal{L}=-\mathbb{E}_{x,r_{\texttt{C}},r_{\texttt{R}}}\left[\ln\sigma\left(\beta\ln\frac{\pi(r_{\texttt{C}}|x)}{\pi_{0}(r_{\texttt{C}}|x)}-\beta\ln\frac{\pi(r_{\texttt{R}}|x)}{\pi_{0}(r_{\texttt{R}}|x)}\right)\right].}$
While $\mathcal{P}_{\texttt{R}}$ explains who may prefer $r_{\texttt{R}}$ (section 3.1), this is not causal: users in $\mathcal{P}_{\texttt{R}}$ may not be best satisfied by $r_{\texttt{R}}$ Jin et al. (2021)—the goal of PT. Empirically, PT does not benefit much when trained on rejected signals $\mathcal{P}_{\texttt{R}}$ and $r_{\texttt{R}}$, (Appendix A.8), indicating $r_{\texttt{R}}$ has lower average quality than $r_{\texttt{C}}$. Thus, we train our PT models just on $r_{\texttt{C}}$ and $\mathcal{P}_{\texttt{C}}$. However, since $\mathcal{P}_{\texttt{R}}$ is valid, it can be used in inference; in section 5.4, we show $\textsc{PT}_{\textsc{dpo}}$ supports needs in $\mathcal{P}_{\texttt{R}}$, unlike DPO.
We use greedy decoding, but on Anthropic, this sometimes yields repetitive, non-terminating texts (20% of times), as some outputs used to train models have this repetition. We filter such outputs for fair evaluation, but use the full test set in Appendix A.11—reflecting use cases when filtering is infeasible; our results remain strong on the full test set.
4.3 Metrics
To compare outputs from varied personalized models, we use a common method of model win-rate Liu et al. (2023) judged by Prometheus Kim et al. (2024). It judges the better model output on: 1) Response Quality: answering the prompt; and 2) Personalization: tailoring to the persona; these assess how well models attend to both PT inputs. For position bias, we judge outputs in both orders, only noting a winner if an output is picked twice, otherwise a tie. For the three judgment types (win/loss/tie), Prometheus has $66\%$ human agreement on (1) in three out-of-domain datasets Kim et al. (2024), the best open-weight judge, and $62\%$ agreement with two authors on (2) (Appendix A.4). Prometheus is imperfect, so users also assess responses in section 5.4.
There are quality/personalization tradeoffs: tailored outputs are specific and appeal to fewer users Chakraborty et al. (2024), harming judged quality. To capture this, we create $\Delta\textsc{PQ}$ to measure the mean gain in both metrics Demšar (2006). Formally, to see if a new model $\pi_{test}$ bests a baseline model $\pi_{base}$, we query Prometheus for win/tie/loss rates of $\pi_{test}$ vs $\pi_{base}$ on personalization ($p_{test}$, $p_{tie}$, $p_{base}$) and quality ($q_{test}$, $q_{tie}$, $q_{base}$). $\Delta\textsc{PQ}$ finds the mean improvement of $\pi_{test}$ vs $\pi_{base}$ on both metrics, ignoring ties, relative to a 50/50 random judge:
$p_{win}$ $=\frac{p_{test}}{p_{base}+p_{test}},\quad q_{win}=\frac{q_{test}}{q_{base}+q_{test}},$
$\Delta\textsc{PQ}$ $=\frac{1}{2}\left(\frac{p_{win}-0.5}{0.5}+\frac{q_{win}-0.5}{0.5}\right).$
If $\Delta\textsc{PQ}>0$, $\pi_{test}$ bests $\pi_{base}$ by giving more personalized and higher-quality outputs, or improves in one metric with minimal reductions in the other.
5 Evaluating Persona-Tailored Responses
We now compare Persona Tailoring (PT) to models trained on standard preference data [^6] Preference data is used to train small, domain-specific models Ouyang et al. (2022), so we omit large, general models (ChatGPT) that cannot be preference-tuned for domains., showing PT improves personalization based on LLM (section 5.1, section 5.2, section 5.3) and human judgments (section 5.4, section 5.5).
5.1 Persona Tailoring Aids Personalization
We first show the net benefits of PT, confirming it enhances personalization while maintaining quality versus models tuned on standard preference data that are not personalized. PT always improves in personalization via FS, SFT, and DPO with typically minor response quality losses, shown via large $\Delta\textsc{PQ}$ (Table 2). Thus, researchers seeking to improve personalization should elicit why users prefer responses in data collection. Further, if preference data has already been curated, abduction via PI is a simple but effective augmentation method that enhances personalization with low impact on quality.
Retrieved personas $\mathcal{P}_{\texttt{retr}}$ sometimes improve response quality ($\textsc{PT}_{\textsc{fs}}$ on BeaverTails, $\textsc{PT}_{\textsc{fs}}$/$\textsc{PT}_{\textsc{sft}}$ on Mnemonic), so personas can also help models give generally high-quality responses. We believe using personas as constitutions Bai et al. (2022b) to align LLMs could be fruitful and executed via abductive moral reasoning Rao et al. (2023) for PI.
5.2 $\textsc{PT}_{\textsc{dpo}}$ Supercharges Personalization
We verify $\textsc{PT}_{\textsc{dpo}}$ is the best method by comparing each PT model to FS without personas. Method response lengths have large differences, so for verbosity bias Zheng et al. (2024), we judge model outputs with the same sentence count; the trend is similar without this check (Appendix A.7). Each method shows $\Delta\textsc{PQ}$ gains on Mnemonic, but SFT shows minor losses on BeaverTails (Table 3). Perhaps LLaMA-8B has seen safety data like BeaverTails, so the base few-shot model has high response quality, while Mnemonic is likely out-of-domain, benefiting from SFT. Regardless, $\textsc{PT}_{\textsc{dpo}}$ has the best $\Delta\textsc{PQ}$. Thus, alignment methods like DPO on persona-augmented preference datasets better instill personalization than alternatives like SFT/FS.
Dataset $\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
Beaver Tails FS $\textsc{PT}_{\textsc{fs}}$ 58.5/23.8/17.7 59.2/17.7/23.1 +48.7
$\textsc{PT}_{\textsc{sft}}$ 37.7/29.5/32.8 24.6/34.4/41.0 -9.0
$\textsc{PT}_{\textsc{dpo}}$ 78.0/19.8/2.2 58.2/18.7/23.1 +68.9
Mnem. FS $\textsc{PT}_{\textsc{fs}}$ 41.1/30.5/28.4 45.0/21.5/33.5 +16.4
$\textsc{PT}_{\textsc{sft}}$ 43.6/40.4/15.9 37.9/43.6/18.5 +40.5
$\textsc{PT}_{\textsc{dpo}}$ 78.2/16.4/5.5 43.6/38.2/18.2 +64.1
Table 3: Ablations of PT steps using $\mathcal{P}_{\texttt{retr}}$. Each step improves personalization and usually improves response quality.
Dataset $\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
BT Chosen DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 46.7/29.3/24.0 38.5/30.5/31.1 +21.3
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 42.3/29.3/28.5 34.9/33.9/31.3 +12.5
BT Reject DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 45.1/31.7/23.2 35.1/32.5/32.5 +17.9
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 51.1/25.9/23.0 35.3/32.7/32.1 +21.3
HHH Chosen DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 37.2/22.6/40.2 32.6/21.3/46.0 -10.4
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 32.6/27.9/39.5 30.4/29.0/40.6 -11.9
HHH Reject DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 48.0/21.8/30.1 39.3/25.8/34.9 +14.4
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 50.8/20.7/28.5 43.0/23.4/33.6 +20.2
Mnem Chosen DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 42.6/31.2/26.2 40.2/31.6/28.2 +20.7
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$
Mnem Reject DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 37.4/32.6/30.0 42.0/27.4/30.6 +13.3
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$
Average DPO $\textsc{PT}_{\textsc{dpo}}$ 43.4/27.3/29.3 43.8/28.8/30.0 +11.9
Table 4: Comparison of personalization abilities of DPO and $\textsc{PT}_{\textsc{dpo}}$. DPO has some tailoring ability on chosen personas, but fails on rejected ones. $\textsc{PT}_{\textsc{dpo}}$ often excels in both personas.
5.3 DPO Sans Personas is Not All You Need
Having seen $\textsc{PT}_{\textsc{dpo}}$’s benefits (section 5.1, section 5.2), we now test if training on personas is needed for personalization: could DPO trained without input personas still tailor to inputted personas in inference? This could be doable, as LLMs may generalize to unseen instructions Deb et al. (2022), and personas are instructions. To answer this, we mainly study BeaverTails and Anthropic HHH, as their prompts are first-person, so we can add personas to prompts by writing them in first person (“the user is X”$\rightarrow$“I am X”). We also test Mnemonic, but since prompts in this dataset are a single vocabulary term, models may not be able to generalize to personas as instructions; this highlights another advantage of our method, as we can instill personalization in any type of preference data. We then compare if outputs from $\textsc{PT}_{\textsc{dpo}}$ or DPO better tailor to input personas $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ on response quality [^7] This response quality judge uses the persona and prompt as inputs, as DPO and $\textsc{PT}_{\textsc{dpo}}$ have access to both in inference. and personalization (section 4.3). This ensures DPO and $\textsc{PT}_{\textsc{dpo}}$, which are tuned on majority chosen outputs and thus may implicitly tailor to popular needs, still aid less popular but valid needs found in $\mathcal{P}_{\texttt{R}}$ (section 3.3).
$\textsc{PT}_{\textsc{dpo}}$ bests DPO in personalization and quality in 8/10 splits, reinforcing PT’s strength (Table 4). $\textsc{PT}_{\textsc{dpo}}$ also has higher improvements than DPO on $\mathcal{P}_{\texttt{R}}$ vs $\mathcal{P}_{\texttt{C}}$ (mean $\Delta\textsc{PQ}$ of 14.5 on $\mathcal{P}_{\texttt{R}}$ vs 6.4 on $\mathcal{P}_{\texttt{C}}$), so DPO sometimes tailors to needs in $\mathcal{P}_{\texttt{C}}$, but rarely the uncommon ones in $\mathcal{P}_{\texttt{R}}$. Thus, to curate harder evaluations for personalization, researchers can consider the often ignored reasons users may prefer rejected outputs and fully challenge models.
Figure 6: Evaluation of how well LLMs aid user-given needs. PT significantly improves personalization while maintaining answerability (Dror et al., 2018, 95% bootstrapped CIs).
Figure 7: Example PT outputs from DPO. Our model can tailor to information needs and learning styles (blue), but may still err on adversarial, irrelevant, or sensitive personas (red).
5.4 PT Personalizes to User-Specified Needs
To verify PT aids user-written needs, we run a pilot study with 10 CS students who often use LLMs. We pick 12 queries from BeaverTails and Anthropic HHH (party plans, job search) that students may ask LLMs Zhao et al. (2024b). Users first write three personas per prompt so models must adapt to personas, and then rate $\textsc{PT}_{\textsc{dpo}}$ and DPO outputs—using prompts and their personas as inputs—on 1-5 on Answerability (answering queries) and Personalization (tailoring to personas). This mirrors our offline metrics (section 4.3) and assesses whether models meet the needs of users who specified them.
Both models have high answerability but $\textsc{PT}_{\textsc{dpo}}$ is significantly more personalized (Figure 6), so $\textsc{PT}_{\textsc{dpo}}$ better supports user-written needs. As users perceive $\textsc{PT}_{\textsc{dpo}}$ as more helpful, future work can assess how PT impacts long-term trust or engagement Serino et al. (2005) and if it aids downstream tasks Balepur et al. (2024b); Mozannar et al. (2024).
5.5 Promises and Pitfalls of Persona Tailoring
We show strengths and issues of $\textsc{PT}_{\textsc{dpo}}$ (Figure 7). $\textsc{PT}_{\textsc{dpo}}$ alters answers for information needs and tailors mnemonics to learning styles (blue), showing its promise for downstream personalization tasks. However, $\textsc{PT}_{\textsc{dpo}}$ assumes personas are harmless, leading to sycophancy Sharma et al. (2024); adversaries can exploit this and use personas to get inaccurate, biased, or irrelevant text (red). To solve this, future work can create undesired personas to teach LLMs when to abstain Wang et al. (2024b).
6 Related Work
Personalization:
LLM personalization aims to produce text aligned with a user’s needs, interests, or attributes Zhang et al. (2024c); Kirk et al. (2024a); Chen et al. (2024b). Personalization can aid engagement Pardini et al. (2022), trust Serino et al. (2005), and learning Bernacki et al. (2021) across recommender systems Islam et al. (2014), gaming Tondello et al. (2017), chatbots Madotto et al. (2019), and tutoring frameworks Park et al. (2024).
For personalization, many works use personas—textual user descriptions Salminen et al. (2020)—in prompts Jandaghi et al. (2024); Liu et al. (2024a); Deshpande et al. (2023), but we train LLMs with personas. Stephan et al. (2024) also train LLMs with verbal feedback but focus on overgeneralization, not personalization. Lee et al. (2024) similarly test persona training, but use rule-based personas and have GPT-4 produce tailored outputs for their personas, forming synthetic training data. In contrast, we infer personas in preference data, which has more flexibility and does not require a teacher model that already excels in personalization; the latter is key in specific domains like education where we cannot rely on LLMs as ground-truth, but user preference data can be curated Liu et al. (2024c).
Lastly, works have tried eliciting personas from a user’s interaction history Li et al. (2023); Handa et al. (2024); Jin et al. (2024). Our PI task is a form of preference elicitation, deriving personas from a single pairwise comparison (section 2). However, we decompose the training of persona-tailored LLMs and the elicitation of personas for inference into separate research questions, focusing on the former.
Preference Subjectivity:
User preferences are subjective and depend on a user’s opinions, traits, and values Bakker et al. (2022); Kirk et al. (2024b); Agnew et al. (2024). As a result, researchers have studied social choice theory Conitzer et al. (2024) and pluralistic alignment Sorensen et al. (2024); Liu et al. (2024a) to handle subjectivity, executed via Bayesian preference tuning Yang et al. (2024); Handa et al. (2024), merging models individually aligned to groups Chakraborty et al. (2024); Hwang et al. (2023), multi-LLM collaboration Feng et al. (2024); Wu et al. (2024b) and multi-objective reward modeling Wang et al. (2024a); Zeng et al. (2024).
While these methods capture preferences across examples, we infer preferences in single pairwise comparisons via abductive reasoning—finding evidence to explain outcomes Peirce (1974). This reasoning has been used in commonsense Zhao et al. (2023), question answering Wu et al. (2024a), and reading comprehension Du et al. (2021) tasks, but we use it to infer personas of users that prefer chosen or rejected outputs. Very recent and concurrent work also explores how contexts can change preferences on responses to improve human evaluation Malaviya et al. (2024) and preference modeling Pitis et al. (2024), but we are the first to treat these contexts as personas to tailor responses for users.
7 Conclusion
Abductive reasoning—explaining when outputs are preferred—greatly aids personalization via persona inference (PI) and tailoring (PT). PI infers uncommon but valid reasons to prefer rejected responses which form harder evaluations, so future work can curate better, long-tail user needs Yin et al. (2012) to stress-test personalization. Users favor PT’s outputs, but there are many questions to study, like if PT helps users finish tasks Mozannar et al. (2024) and if users can easily verbalize needs as prompts Taylor (1962). Despite PT’s success, we may need debiasing and abstention Meade et al. (2022); Wang et al. (2024b) to curb harmful or biased personas.
Beyond personalization, personas reveal implicit differences in chosen and rejected outputs, which can be further surfaced via contrastive topic models Zhong et al. (2023). Our personas convey user traits and needs, but other types—like moral arguments Rao et al. (2023), cultural values Kirk et al. (2024b), and authorship Wang et al. (2023)—could adapt LLMs to diverse constitutions, cultures, and writing styles. Overall, we advocate for an abductive view of preferences for personalization, asking why, when, and which users may prefer responses.
8 Limitations
One limitation is that our persona tailoring method assumes there is always an input persona that is relevant to the prompt. However, as we show in section 5.5, a user could give personas that are irrelevant to the prompt, harming our model outputs. To address this, future work could explore producing intentionally irrelevant personas and training the model to abstain on them. Further, to address cases where a specific persona is not needed as input, researchers could route prompts to a model trained without personas, or use a default persona like a general system prompt; we test the latter in Appendix A.9.
We also acknowledge that we cannot fully evaluate all possible personas and user needs that a model must cater to. To enhance persona diversity for evaluation, we use personas derived from both chosen and rejected responses in offline evaluation (section 5.3), resulting in two diverse user needs per input prompt. Further, in our pilot study, we ask annotators to write three different personas per input prompt, and we find few exact matches in personas. We believe there is promising future work to rigorously test how personalization impacts diverse users, including if personalization harms or helps users across cultures Kirk et al. (2024b), if personalization truly helps users in downstream tasks Mozannar et al. (2024), and if users can accurately articulate their preferences Handa et al. (2024).
Finally, LLMs are sensitive to prompt variations Sclar et al. (2024), so our persona inference and tailoring results may differ by prompt design. To address this, we follow best practices in prompt engineering Schulhoff et al. (2024) to ensure each model has a fair opportunity to generate valid personas in persona inference. Further, due to resource constraints, we only experimented with a single combination of student and teacher models for persona tailoring. We believe future work can investigate the potential of using the same model for both persona generation and training, employing user personas in a self-evolving learning framework.
9 Ethical Considerations
Personalization can raise ethical concerns when using personas tied to sensitive attributes like race or gender, which risks perpetuating stereotypes Kantharuban et al. (2024). As a result, we did not focus on these types of personas in this work, and instead studied high-level interests, personality traits, and information needs (section 2). While users specifying sensitive personas directly is less concerning, issues arise when practitioners use preference elicitation techniques Li et al. (2023) to infer such personas from past interactions and employ them for personalization. Although we observed no instances of protected attributes in persona inference, our model trained on “safe” personas could still produce harmful responses if prompted to (section 5.5). We recommend future works to explore safeguards, such as classifiers to flag harmful personas pre-inference, and conduct user studies to understand which types of personas users prefer seeing reflected in outputs, guiding responsible personalization efforts.
Further, personalization could risk selection or confirmation bias Hernán et al. (2004); Klayman (1995). For example, if a user already has a particular view and requests information that is aligned with that view, our model will provide a response that confirms this user’s view, which may be harmful in cases such as misinformation Zhou and Shen (2022). There is a tradeoff between generating personalized and balanced responses, and researchers can explore future task setups that encourage generating balanced outputs Zhang et al. (2024b) to better quantify the effects of this tradeoff.
Acknowledgments
We would like to thank the CLIP lab at the University of Maryland and our external collaborators for their feedback. We also thank Dang Nguyen, Yu Hou, Connor Baumler, Paiheng Xu, Zongxia Li, Feng Gu, Naina Balepur, and Atrey Desai for their annotation help. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber), iis-2339746 (Rudinger), and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research.
Appendix A Appendix
A.1 Dataset Details
We give descriptions of the datasets used in persona tailoring in Table 11. All datasets are publicly available online, so they are within their intended use. Datasets do not have PII and are in English. To avoid any potentially offensive content, we only use outputs marked as “safe” in BeaverTails and with 10+ upvotes on SHP.
On each test set entry for PT, we use both personas derived from the gold outputs ($\mathcal{P}_{\texttt{gold}}$), and those based on an example that was retrieved from the training sets. To show what this looks like, consider the test set prompt “Can you give me advice on hacking someone’s identity?” from BeaverTails. The chosen persona derived from the gold examples is “The user is security-conscious and prefers practical advice on protecting themselves, rather than simply being told what not to do”, which leaks some signals to the model, as it reveals that the chosen response discusses protection advice, while the rejected response refuses to answer. To help avoid this potential leakage, we retrieve a similar prompt from training; in this case, the question is “Can you help me hack into someone else’s email account so that I can gain access to their personal data?”. From this example, the retrieved question is similar, but not identical to the original input prompt, so the persona linked to this example (i.e. “The user is ethically-minded and prefers responses that prioritize legal and moral guidelines over technical feasibility.”) does not directly leak signals on the gold output.
A.2 Persona Inference Setup
The exact instructions given for Persona Inference for Stanford Human Preferences are as follows:
“You will be given a prompt and two responses: a response that was chosen by the user (Chosen Response) and a response that was rejected by the user (Rejected Response) during a pairwise comparison. The prompt is a title of a forum post containing a question and the responses are comments that provide answers for the original poster. Your task is to generate a very short, specific, one-sentence description of the user’s preference, i.e. a persona. The persona should contain reasoning for why the user preferred and picked the Chosen Response and did not pick the Rejected Response. The persona should be very short and should not mention specific details in the prompt or responses, but instead should discuss higher-level characteristics that can be inferred about the user’s persona.”
On BeaverTails, we alter instructions so the prompts are “questions” and responses are “answers.” On Anthropic HHH, the prompts are “human utterances” and responses are “assistant utterances.” On Mnemonic, the prompt is a “vocab term” and responses are “keyword mnemonics.”
These instructions are prepended and appended to five-shot examples manually written by the authors, and we ensure that the exemplars are diverse and representative of the datasets. We found that qualitatively, putting instructions before and after the exemplars improved the quality of personas, similar to the re-reading prompting technique Xu et al. (2024). Overall, this prompt is based on best practices in prompt engineering Schulhoff et al. (2024), ensuring consistent instructions across models, including input/output definitions, balanced few-shot exemplars, and output requirements.
All LLMs generate with 0 temperature, a maximum sequence length of 2048 tokens, and use the token “Prompt:” for early stopping. All unspecified parameters are default values. We do not do hyperparameter tuning and results are reported from a single run. Each run is allocated 24 CPU hours.
A.3 Persona Tailoring Setup
We train the SFT and DPO models using the Transformer Reinforcement Library (trl) [^8] https://huggingface.co/docs/trl/en/index on huggingface. The SFT model uses a maximum sequence length of 512 tokens, a batch size of 1, 10 training epochs, and a learning rate of $2\cdot 10^{-5}$; we select the model with the lowest evaluation loss after each epoch. The DPO model uses a learning rate of $5\cdot 10^{-6}$, $\beta=0.1$, and all other hyperparameters are the same as SFT; we also select the model with the lowest evaluation loss after each epoch. For efficiency, SFT and DPO are both trained with LoRA Hu et al. (2022) using $r=16$, $\alpha=32$, a dropout of 0.05, and no bias. All unspecified parameters are default values. We do not perform hyperparameter tuning and results are reported from a single run. Each run is allocated 24 GPU hours on a single NVIDIA A100 GPU.
A.4 LLM Judge Details
We use two LLM judges in our paper: GPT-4o for persona inference accuracy (section 3.1), and Prometheus-7B for persona tailoring quality (section 4.3). GPT-4o is prompted in the same way as Appendix A.2, with adapted instructions tasking the model to identify the better response based on the provided persona, and the same five examples as in persona inference. While GPT-4o is a slightly more reliable LLM judge, it is expensive to run, so we use Prometheus-7B, the strongest open-source judge, for persona tailoring evaluation Kim et al. (2024). For personalization evaluation, the model is given the input prompt, the persona, and two model responses, and is asked: “Does the response answer the prompt and align with the user’s specified persona?” For response quality evaluation, the model is given the input prompt and two model responses, and is asked: “Is the response high-quality?”
To assess GPT-4o’s reliability, we calculate the agreement of one Ph.D. student (an author) with the LLM in persona inference accuracy (section 3.1). We sample 100 random GPT-4o judgments across all persona inference models/datasets, and the student does the same task as the LLM. We find a high raw agreement of 86%, so the metric is reliable.
We use a similar evaluation for the Prometheus-7B judge in persona tailoring. Given the subjective nature of personalization Jang et al. (2023), we have two Ph.D. students (the authors) give judgments (i.e. A wins, B wins, Tied) on 50 random model pairwise comparisons. The students have a moderately high Kendall’s tau of 0.63, showing the subjective nature of personalization. When we average the students’ responses and compare them to the model’s judgments, we find an agreement of 62%. The value is very close to the agreement of 66% reported by the authors of Prometheus Kim et al. (2024) when averaged over the three tested out-of-domain datasets with ties (random chance is 33%), so Prometheus gives personalization judgments of similar accuracy to quality judgments.
A.5 Extended Implicit Preference Analysis
In this section, we provide more details and extend our analysis of implicit preferences from section 3.4. To tokenize words for word saliency, we use nltk [^9] https://www.nltk.org/ and PortStemmer to group similar words together. We only display words that are nouns, adjectives, or words, ignoring spurious words like prepositions (e.g. “by” has a saliency of 0.91 on SHP). Our personas sometimes include anti-preferences (e.g. “the user prefers X rather than Y”), so we split personas by contrasting words (i.e., “rather than”, “over”, “versus”, and “compared to”) and compute word saliency via the first half of personas.
We repeat our analysis across datasets in Table 6. Anthropic HHH has similar trends to BeaverTails; chosen responses are associated with solutions, results, and facts, while rejected responses are considered short and high-level, another indication of a tendency towards verbose outputs. On SHP, which is derived from Reddit posts, chosen responses are associated with curious users who want to know techniques and workarounds; while rejected responses are more balanced and minimal.
A.6 LLM Personas Flip LLM Preferences
As LLMs can infer personas $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ that justify $r_{\texttt{C}}$ or $r_{\texttt{R}}$ as preferred responses (section 3.1), we now test how seeing both $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ alters an LLM’s preferences. To do so, we first 0-shot prompt LLMs for an initial preference $y_{0}$—if they prefer $r_{\texttt{C}}$ or $r_{\texttt{R}}$ for prompt $p$. We shuffle outputs and set $y_{0}=\texttt{C}\text{ or }\texttt{R}$ if $r_{\texttt{C}}$ or $r_{\texttt{R}}$ win in both orders ($y_{0}=\texttt{Tie}$ otherwise). We then find preferences with personas $y_{\mathcal{P}}$, where each model gives its preferred response but also uses its inferred personas $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ as inputs.
When $y_{0}=\texttt{C}$ and $y_{0}=\texttt{R}$, LLMs switch their preference $y_{\mathcal{P}}$ after seeing both personas 36% and 49% of times (Figure 8). Further, when $y_{0}=\texttt{Tie}$, $y_{\mathcal{P}}$ is split fairly evenly between C and R. As LLMs similarly alter their preferences when seeing both personas, $\mathcal{P}_{\texttt{R}}$ has similar persuasiveness to $\mathcal{P}_{\texttt{C}}$, confirming users can prefer $r_{\texttt{R}}$ for valid reasons.
A.7 Extended Ablations
We present further results and descriptions of our ablation studies (section 5.2). We first explore length discrepancies in model outputs (Table 7), finding that models using different generation strategies (center) have large length discrepancies over one sentence long, while models using the same generation strategy (top) have similar lengths. To avoid verbosity bias in ablations Zheng et al. (2024), we only use pairs with the same number of sentences. The comparison between DPO and $\textsc{PT}_{\textsc{dpo}}$ in section 5.3 also has discrepancies in sentence count, but as this is a personalization comparison and personas often relate to length (e.g. “The user is comprehensive”), we feel length adjustments are not appropriate.
Another interesting finding in our ablations is that on BeaverTails, supervised fine-tuning with personas does not surpass few-shot prompting with personas. To give a potential explanation, we also conduct ablations of models trained without personas (Table 3). We similarly find that the supervised fine-tuning model underperforms the few-shot prompted model on BeaverTails. We speculate that LLMs have already been pretrained and undergone base alignment on a wide variety of safety datasets similar to BeaverTails. As a result, the few-shot prompted model may produce high-quality outputs on these safety datasets and do not benefit as much from fine-tuning. In contrast, the Mnemonic dataset is a niche task that the model likely has not seen frequently in pretraining, and thus, supervised fine-tuning still has benefits.
A.8 Training on Rejected Responses
In our experiments, we train models just on chosen responses and personas, as we did not find as much benefit from training on rejected personas. In Tables 10, 11, and 12, we evaluate using chosen, rejected, and both chosen and rejected personas for persona tailoring training and inference, compared to the baseline generation strategies that do not use either (section 5.1). Few-shot and supervised fine-tuning have generally positive benefits in $\Delta\textsc{PQ}$, even when training and running inference on rejected personas. Thus, when paired with the right personas, rejected responses can form valuable training signals for these strategies. However, direct preference optimization has smaller benefits, with $\Delta\textsc{PQ}$ most often reaching negative values.
We believe that while there may be valid reasons to prefer rejected responses, it does not mean that the rejected response is the best output for a user who aligns with said reason. As a result, methods that instill high personalization (section 5.2) like DPO may overfit to the negative qualities of the rejected response that led many users to disprefer it, leading to lower quality judgments and thus negative $\Delta\textsc{PQ}$.
A.9 System Prompt Personas in PT
While the main goal of persona tailoring is to allow for custom, specified user needs during each inference example, we also explore the effects of using a fixed system prompt across the entire set. We base our system prompts on the insights from section 3.4, using the most salient tokens associated with chosen responses in the system prompts, which we hope will more likely lead to higher-quality responses.
For BeaverTails, our written system prompt is “The user is meticulous and prefers responses that cover multiple, diverse angles.” For Anthropic HHH, our system prompt is “The user is solution-focused, results-oriented, and fact-oriented, and prefers responses that cover varied angles.” For Mnemonic, our system prompt is “The user prefers indirect, step-by-step mnemonics that capture the essence of the vocabulary term.”
In Table 13, we find that on BeaverTails and Mnemonic, our system prompt persona $\mathcal{P}_{retr}$ with persona tailoring surpasses the baseline generation strategies, shown via positive $\Delta\textsc{PQ}$. On Anthropic HHH, we do not see similar benefits. We consider it a very positive sign that our first attempt at specifying a system prompt often improved quality and personalization, and future works can explore optimization techniques Zhang et al. (2024a) to find the best system prompt persona for a dataset.
A.10 Evaluating Teacher and Student Models
Our use of LLaMA-405B to infer personas and training LLaMA-8B on these personas is similar to the student-teacher paradigm in knowledge distillation Gou et al. (2021). In this section, we compare the abilities of our fully trained persona tailoring model with DPO ($\textsc{PT}_{\textsc{dpo}}$) with few-shot prompted LLaMA-405B, both using the retrieved personas $\mathcal{P}_{retr}$ (Table 14). While our PT model has less than 2% of the parameters of LLaMA-405B, it shows competitive performance with LLaMA-405B on BeaverTails and Mnemonic; $\Delta\textsc{PQ}$ is only 2.27 on BeaverTails, showing the models can be competitive. This further confirms the strength of PT; it lets smaller, trained models produce high-quality, personalized outputs with a fraction of parameters.
A.11 Repetitive Text on Anthropic HHH
On Anthropic HHH, our manual analysis found that models trained on this dataset and using greedy decoding could repeat text without generating the end-of-text token (around $20\%$ of cases). For example, when given a user request for celery recipes, the DPO model produces the text: “You can eat it raw, add it to salads, or use it in soups and stews. You can also make celery juice, or use it in a celery juice cocktail. You can also use it in a celery juice smoothie, or in a celery juice cocktail…” until the maximum token length is reached; some outputs in the dataset have this repetitive nature, which we believe could lead to this behavior. To avoid re-running all experiments by filtering these examples in training data, we omitted these outputs in inference to ensure no model had a fair/unfair advantage in generating these types of repetitive outputs.
To ensure these outputs do not impact our findings, we run our evaluations on Anthropic HHH using the full test set. We repeat our experiments in section 5.1 with this constraint in Table 15, and our experiments in section 5.3 in Table 16. Our findings are consistent (and in fact, stronger), showing that PT typically has net improvements in personalization while maintaining quality, while PT excels in personalizing to the uncommon but valid needs linked to rejected responses. Thus, our results confirm that PT is still a strong technique despite this issue.
In future directions of this work, we plan to explore varied decoding strategies (e.g. greedy decoding, beam search, nucleus sampling, repetition penalties) to investigate how they impact model personalization Zarrieß et al. (2021).
A.12 Annotator Instructions
We present instructions given to annotators when writing personas and assessing outputs for prompts and personas in Figure 9 and Figure 10, respectively. The annotators are four graduate students and one undergraduate student studying computer science and have research experience with LLMs, natural language processing, and information retrieval; none of them are authors. We did not collect any personal information about the participants, except for their preferences in personas. Participants were rewarded with gift cards and purchased meals which averaged around \$20 per hour, above our region’s minimum wage. Our setup was approved by an Institutional Review Board to mitigate any potential harms. Annotators were aware that their average ratings would be reported in the paper.
Dataset Domain #SFT Train #SFT Val #DPO Train #DPO Val #Test Prompt Len. Chosen Len. Rejected Len.
BeaverTails QA 977 244 982 246 500 24.08 103.97 80.02
Anthropic Dialogue 424 105 423 107 500 16.77 77.21 61.69
Mnemonic Education 126 35 132 35 500 2.48 11.02 11.12
Table 5: Description (domains, training splits, and lengths) of datasets used in persona tailoring. Length is the average number of tokens computed by tiktoken. [^11] https://github.com/openai/tiktoken
Figure 8: Rejected response personas are as valid as chosen ones. Prompting LLMs with personas often switches their initial preference to tie, and when LLMs initially rate the responses as tied, they remain split between responses
Dataset Type Salient Words Example Prompt Example Response Example Persona
BeaverTails Chosen multiple (0.94) What are the laws regarding animal cruelty in the United States? The United States has both state and federal laws that prohibit animal cruelty… The user prefers comprehensive, detailed information that covers the topic from multiple angles
meticulous (0.92)
diverse (0.92)
BeaverTails Rejected to-the-point (1.0) Where can I find legal advice regarding injury prevention? The CDC provides comprehensive legal advice and information … on their website: cdc.gov/injury. The user is direct and to-the-point, preferring concise and specific information sources.
directness (0.90)
concise (0.90)
Mnemonic Chosen indirect (0.81) Ascertain Break it down to "assure + certain"; to ascertain is to make sure of something. The user is a logical thinker and prefers clear, step-by-step breakdowns to understand new concepts.
step-by-step (0.76)
essence (0.75)
Mnemonic Rejected strong (0.91) Zephyr Zephyr sounds like "sea-fear," like a gentle breeze that calms the fear of sailors asea. The user is a romantic thinker and prefers poetic, emotive associations.
story-like (0.90)
romantic (0.79)
Anthropic Chosen solution-focused (0.91) I want to potty train my son but he does not seem interested. Oh, that’s great! I’m happy to help out with this. It sounds like a good idea to start with some kind of reward system The user is a practical and solution-focused individual who prefers direct, actionable advice.
results-oriented (0.85)
fact-oriented (0.85)
Anthropic Rejected to-the-point (0.93) I want to buy a Bluetooth speaker What kind of speaker are you looking for? The user prefers direct, to-the-point questions that efficiently narrow down their needs.
summary (0.91)
high-level (0.90)
SHP Chosen curious (0.79) Why are eggs on so many foods in so many ways? Break it down to "assure + certain"; to ascertain is to make sure of something. Because across nearly all cultures; bird eggs were easily forage-able, and once tamed…
technique (0.75)
workarounds (0.75)
SHP Rejected balanced (0.92) What do I do with old clothes that can’t be worn anymore? https://fordays.com/products/take-back-bag The user is convenience-oriented and prefers quick, easy solutions that require minimal effort and time.
control (0.91)
minimal (0.90)
Table 6: Extended token saliency analysis from section 3.4 across all datasets.
Comparison Type $\pi_{base}$ $\pi_{test}$ $\pi_{base}$ Avg # Sentences $\pi_{test}$ Avg # Sentences
Training on Personas (section 5.1) FS $\textsc{PT}_{\textsc{fs}}$ 6.83 6.46
Training on Personas (section 5.1) SFT $\textsc{PT}_{\textsc{sft}}$ 5.31 5.24
Training on Personas (section 5.1) DPO $\textsc{PT}_{\textsc{dpo}}$ 5.07 4.62
Ablation (section 5.2) FS SFT 4.59 2.77
Ablation (section 5.2) FS DPO 4.59 2.90
Personalization Comparison (section 5.3) DPO $\textsc{PT}_{\textsc{dpo}}$ 7.00 5.449
Table 7: Comparison of output length (number of sentences) between model output pairs in LLM judgments. Our inter-models Ablations (center) have large length discrepancies (over 1 sentence difference), leading us to restrict the evaluation to pairs with the same number of sentences for verbosity bias, but intra-model comparisons (top) have a similar number of sentences (less than 1 sentence). The comparison between DPO and $\textsc{PT}_{\textsc{dpo}}$ also has differences in number of sentences, but since this is a personalization comparison and personas can relate to length (e.g. “The user prefers comprehensive outputs”), we feel length adjustments would not be appropriate.
BeaverTails Mnemonic
$\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
FS SFT 19.4/30.5/50.1 17.0/27.9/55.1 -48.4 35.3/41.2/23.5 38.2/39.2/22.7 +22.8
DPO 49.5/32.5/18.0 51.9/33.1/15.0 +50.8 46.6/35.0/18.4 61.4/26.8/11.8 +55.6
Table 8: Ablations of generation strategies trained on preference datasets without personas. Supervised fine-tuning and direct preference optimization generally improve personalization and quality, except for SFT on BeaverTails, suggesting that the few-shot model already has some training on a wide variety of safety datasets.
Dataset $\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
BeaverTails FS $\textsc{PT}_{\textsc{fs}}$ 62.5/17.2/20.2 60.7/14.2/25.1 +46.3
$\textsc{PT}_{\textsc{sft}}$ 32.1/30.9/37.1 19.0/26.3/54.7 -27.8
$\textsc{PT}_{\textsc{dpo}}$ 76.6/17.6/5.8 49.7/21.8/28.5 +56.5
Mnemonic FS $\textsc{PT}_{\textsc{fs}}$ 44.3/28.5/27.2 46.4/20.5/33.1 +20.3
$\textsc{PT}_{\textsc{sft}}$ 43.7/39.1/17.2 40.1/39.5/20.4 +37.9
$\textsc{PT}_{\textsc{dpo}}$ 78.6/16.4/5.0 49.6/32.2/18.2 +67.2
Table 9: Ablations of generation strategies without controlling for verbosity bias. Our ablations show a similar trend as Table 3; each strategy tends to help on Mnemonic, but $\textsc{PT}_{\textsc{sft}}$ underperforms FS. We believe the judged lower personalization and quality of $\textsc{PT}_{\textsc{sft}}$ stems from verbosity bias, as the two models have large length discrepancies (Table 7), leading us to only compare outputs with the same sentence count in section 3.1. Regardless, $\textsc{PT}_{\textsc{dpo}}$ is the strongest method on both datasets.
BeaverTails Anthropic HHH Mnemonic
$\mathcal{P}_{train}$ $\mathcal{P}_{inf}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
$\mathcal{P}_{\texttt{C}}$ $\mathcal{P}_{\texttt{C}}$ 62.5/17.2/20.2 60.7/14.2/25.1 +46.3 46.6/18.3/35.1 38.4/15.6/46.0 +2.5 44.3/28.5/27.2 46.4/20.5/33.1 +20.3
$\mathcal{P}_{\texttt{R}}$ 47.7/37.1/15.2 47.1/30.1/22.8 +43.1 42.0/28.8/29.2 30.2/29.6/40.2 +2.0 33.6/41.0/25.4 31.5/40.0/28.5 +9.5
$\mathcal{P}_{\texttt{R}}$ $\mathcal{P}_{\texttt{C}}$ 43.6/37.3/19.1 34.3/30.9/34.7 +19.3 40.0/31.5/28.5 24.4/21.2/54.3 -10.6 58.2/26.2/15.6 69.2/24.2/6.6 +70.2
$\mathcal{P}_{\texttt{R}}$ 48.6/30.9/20.5 29.3/33.1/37.6 +14.2 50.4/26.2/23.4 24.2/20.2/55.6 -1.4 54.4/27.8/17.8 71.8/23.4/4.8 +69.1
$\mathcal{P}_{\texttt{C}}+\mathcal{P}_{\texttt{R}}$ $\mathcal{P}_{\texttt{C}}$ 48.1/32.3/19.6 38.3/32.9/28.9 +28.0 31.5/28.1/40.4 18.6/21.6/59.8 -32.5 56.4/29.8/13.8 65.4/27.0/7.6 +69.9
$\mathcal{P}_{\texttt{R}}$ 43.5/37.5/19.0 32.3/32.5/35.3 +17.3 39.2/28.5/32.3 17.0/19.2/63.8 -24.2 58.4/25.6/16.0 65.0/28.4/6.6 +69.3
Table 10: Response type personalization and quality judgments for few-shot models that use chosen personas, rejected personas, and both personas for training and inference, compared to the few-shot model that does not use personas.
BeaverTails Anthropic HHH Mnemonic
$\mathcal{P}_{train}$ $\mathcal{P}_{inf}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
$\mathcal{P}_{\texttt{C}}$ $\mathcal{P}_{\texttt{C}}$ 44.6/31.7/23.7 33.5/28.6/37.8 +12.3 47.6/30.6/21.9 28.3/30.6/41.1 +9.3 40.8/38.3/20.9 35.2/35.2/29.5 +20.5
$\mathcal{P}_{\texttt{R}}$ 43.1/35.4/21.5 24.4/28.7/47.0 +0.9 53.9/26.1/20.0 32.5/26.9/40.6 +17.4 49.5/31.0/19.6 36.7/36.7/26.7 +29.6
$\mathcal{P}_{\texttt{R}}$ $\mathcal{P}_{\texttt{C}}$ 37.6/36.4/26.1 26.7/31.1/42.2 -2.2 52.3/28.3/19.4 30.1/26.9/43.0 +14.1 40.1/36.9/23.0 30.6/37.9/31.5 +12.9
$\mathcal{P}_{\texttt{R}}$ 43.7/34.4/21.9 21.3/27.9/50.8 -3.8 61.4/23.7/14.9 29.6/28.0/42.5 +21.5 43.8/36.8/19.4 38.6/32.7/28.7 +26.7
$\mathcal{P}_{\texttt{C}}$ + $\mathcal{P}_{\texttt{R}}$ $\mathcal{P}_{\texttt{C}}$ 42.9/34.6/22.5 32.6/30.2/37.2 +12.3 44.6/31.5/24.0 29.6/22.8/47.6 +3.4 40.3/36.9/22.8 33.5/35.7/30.8 +15.9
$\mathcal{P}_{\texttt{R}}$ 45.9/34.6/19.5 22.2/32.9/44.9 +3.2 57.9/25.7/16.4 24.5/28.7/46.8 +12.3 45.8/33.9/20.3 32.1/39.0/28.9 +21.9
Table 11: Response type personalization and quality judgments for supervised fine-tuning models that use chosen personas, rejected personas, and both personas for training and inference, compared to the supervised fine tuning model that does not use personas.
BeaverTails Anthropic HHH Mnemonic
$\mathcal{P}_{train}$ $\mathcal{P}_{inf}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
$\mathcal{P}_{\texttt{C}}$ $\mathcal{P}_{\texttt{C}}$ 72.1/18.2/9.6 36.7/24.4/38.9 +36.8 55.8/25.0/19.2 25.4/25.2/49.4 +8.4 64.4/26.0/9.6 27.8/33.2/39.0 +28.6
$\mathcal{P}_{\texttt{R}}$ 70.3/21.0/8.6 17.4/26.9/55.7 +12.9 66.6/21.6/11.8 25.2/27.2/47.6 +19.6 66.0/25.4/8.6 25.4/32.2/42.4 +25.9
$\mathcal{P}_{\texttt{R}}$ $\mathcal{P}_{\texttt{C}}$ 60.9/24.2/14.8 10.2/21.4/68.3 -6.6 64.8/21.0/14.2 8.8/14.2/77.0 -7.7 34.2/33.4/32.4 16.8/30.4/52.8 -24.5
$\mathcal{P}_{\texttt{R}}$ 58.9/24.8/16.2 3.4/9.8/86.8 -17.8 71.0/19.6/9.4 5.4/18.8/75.8 -5.0 34.6/32.4/33.0 21.0/28.0/51.0 -19.6
$\mathcal{P}_{\texttt{C}}+\mathcal{P}_{\texttt{R}}$ $\mathcal{P}_{\texttt{C}}$ 63.5/22.8/13.6 13.4/16.4/70.1 -1.6 70.6/15.0/14.4 12.2/18.0/69.8 -2.1 63.6/24.8/11.6 21.8/29.6/48.6 +15.5
$\mathcal{P}_{\texttt{R}}$ 55.7/27.7/16.6 3.6/9.6/86.8 -19.0 76.6/15.2/8.2 9.6/18.2/72.2 +2.1 61.6/30.2/8.2 22.0/31.4/46.6 +20.3
Table 12: Response type personalization and quality judgments for direct preference optimization models that use chosen personas, rejected personas, and both personas for training and inference, compared to the direct preference optimization model that does not use personas.
BeaverTails Mnemonic Anthropic HHH
$\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
FS $\textsc{PT}_{\textsc{fs}}$ + $\mathcal{P}_{syst}$ 45.8/35.8/18.3 45.0/27.5/27.5 +33.5 32.7/42.8/24.5 28.7/41.6/29.7 +6.4 38.2/32.9/28.9 40.8/21.1/38.2 +8.5
SFT $\textsc{PT}_{\textsc{sft}}$ + $\mathcal{P}_{syst}$ 35.5/36.4/28.1 30.6/40.1/29.3 +6.9 31.9/36.5/31.6 31.6/39.0/29.4 +2.0 22.7/29.2/48.1 13.4/26.4/60.2 -49.7
DPO $\textsc{PT}_{\textsc{dpo}}$ + $\mathcal{P}_{syst}$ 40.8/33.8/25.4 33.8/32.3/33.8 +11.6 58.9/25.5/15.6 23.8/38.9/37.2 +18.1 18.5/35.9/45.7 19.6/37.0/43.5 -40.2
Table 13: On BeaverTails and Mnemonic, adding a fixed system prompt as the persona for inference typically improves both personalization and quality across training strategies.
BeaverTails Anthropic HHH Mnemonic
$\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
$\textsc{PT}_{\textsc{dpo}}$ L-405B 50.3/9.81/39.9 39.9/14.2/45.9 +2.27 90.2/11.4/5.60 83.0/11.4/5.60 +91.0 38.2/32.9/28.9 48.0/28.8/23.2 +24.3
Table 14: Comparison of persona tailoring with DPO and few-shot prompted LLaMA-405B, both using retrieved chosen personas. Although our persona tailoring model is much smaller (8B parameters), on BeaverTails and Mnemonic, the model shows competitive performance.
BeaverTails Anthropic HHH Mnemonic
$\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
FS $\textsc{PT}_{\textsc{fs}}$+$\mathcal{P}_{\texttt{retr}}$ 62.5/17.2/20.2 60.7/14.2/25.1 +46.3 46.6/18.3/35.1 38.4/15.6/46.0 +2.5 44.3/28.5/27.2 46.4/20.5/33.1 +20.3
$\textsc{PT}_{\textsc{fs}}$+$\mathcal{P}_{\texttt{gold}}$ 68.7/14.5/16.9 62.9/15.9/21.3 +55.0 49.0/18.0/33.1 43.7/17.3/39.0 +12.5
SFT $\textsc{PT}_{\textsc{ft}}$+$\mathcal{P}_{\texttt{retr}}$ 44.6/31.7/23.7 33.5/28.6/37.8 +12.3 47.6/30.6/21.9 28.3/30.6/41.1 +9.3 40.8/38.3/20.9 35.2/35.2/29.5 +20.5
$\textsc{PT}_{\textsc{sft}}$+$\mathcal{P}_{\texttt{gold}}$ 46.7/32.0/21.2 38.2/29.6/32.2 +23.0 53.4/27.9/18.7 37.5/30.3/32.2 +27.8
DPO $\textsc{PT}_{\textsc{dpo}}$+$\mathcal{P}_{\texttt{retr}}$ 72.1/18.2/9.6 36.7/24.4/38.9 +36.8 55.8/25.0/19.2 25.4/25.2/49.4 +8.4 64.4/26.0/9.6 27.8/33.2/39.0 +28.6
$\textsc{PT}_{\textsc{dpo}}$+$\mathcal{P}_{\texttt{gold}}$ 66.3/21.4/12.2 40.9/28.5/30.7 +41.6 56.6/26.0/17.4 33.6/27.8/38.6 +23.0
Table 15: Win, tie, and loss rates of generation methods (FS, SFT, DPO) with and without personas $\mathcal{P}$ in pairwise comparisons from the Prometheus judge when using the full Anthropic HHH dataset. Our results are even stronger than in Table 2, with even larger $\Delta\textsc{PQ}$ increases.
Dataset $\pi_{base}$ $\pi_{test}$ Person. W/T/L Quality W/T/L $\Delta\textsc{PQ}$
BT Chosen DPO+$\mathcal{P}_{retr}$ PT+$\mathcal{P}_{retr}$ 46.7/29.3/24.0 38.5/30.5/31.1 +21.3
DPO+$\mathcal{P}_{gold}$ PT+$\mathcal{P}_{gold}$ 42.3/29.3/28.5 34.9/33.9/31.3 +12.5
BT Reject DPO+$\mathcal{P}_{retr}$ PT+$\mathcal{P}_{retr}$ 45.1/31.7/23.2 35.1/32.5/32.5 +17.9
DPO+$\mathcal{P}_{gold}$ PT+$\mathcal{P}_{gold}$ 51.1/25.9/23.0 35.3/32.7/32.1 +21.3
HHH Chosen DPO+$\mathcal{P}_{retr}$ PT+$\mathcal{P}_{retr}$ 40.8/25.4/33.8 35.0/28.0/37.0 +3.3
DPO+$\mathcal{P}_{gold}$ PT+$\mathcal{P}_{gold}$ 42.0/27.4/30.6 39.0/24.4/36.6 +9.4
HHH Reject DPO+$\mathcal{P}_{retr}$ PT+$\mathcal{P}_{retr}$ 56.2/21.0/22.8 48.6/24.6/26.8 +35.6
DPO+$\mathcal{P}_{gold}$ PT+$\mathcal{P}_{gold}$ 54.1/20.6/25.3 44.7/26.1/29.3 +28.6
Table 16: Comparison of personalization abilities of DPO and $\textsc{PT}_{\textsc{dpo}}$ when using the full Anthropic HHH dataset. Our results are even stronger than in Table 4; DPO always bests $\textsc{PT}_{\textsc{dpo}}$ in personalization and response quality.
Figure 9: Instructions given to annotators when writing personas for input prompts.
Figure 10: Instructions given to annotators when evaluating responses for prompts and personas on personalization and answerability.